# Enabling Training-Free Semantic Communication Systems with Generative Diffusion Models

### v1

@@ Abstract
Semantic communication, joint source-channel coding, diffusion models, image transmission.
@@ I Introduction
To eliminate the dependency on training, recent studies have explored training-free generative SemCom frameworks. Specifically, the authors in [12] introduced a channel-aware GAN inversion method for semantic encoding and employed the same GAN generator for decoding, thus avoiding any encoder–decoder training on the communication task. Meanwhile, the works in [13, 14] proposed to use a pretrained image captioner to extract textual descriptions from the transmitted image. These captions, along with auxiliary visual features such as edge maps or latent vectors, are transmitted to the receiver, where the caption serves as a conditioning prompt to guide a generative diffusion model in reconstructing the original image. However, these training-free approaches lack robustness to channel noise, as the transmitted semantic conditions, such as captions or latent features, are easily corrupted during transmission, leading to degraded reconstruction quality.
To overcome these limitations, we propose a fully training-free generative SemCom framework that leverages publicly available pretrained generative diffusion models (GDMs). Specifically, we design a semantic encoding and decoding method based on the inversion and sampling process of the denoising diffusion implicit model (DDIM)[9], which introduces a two-stage forward diffusion strategy split between the transmitter and receiver to enhance robustness against channel noise. In addition, we optimize the number of sampling steps at the receiver to match the total noise level introduced by the channel noise. We further provide a brief analysis of this design, how channel noise shifts the latent distribution away from the latent distribution of GDM in ideal conditions, and give insights into how the proposed system mitigates this mismatch. Simulations on the Kodak dataset are conducted to demonstrate the superiority of the proposed system across a wide range of perceptual and distortion metrics. In particular, our method achieves over 50% performance gains in Fréchet Inception distance (FID) compared to baselines when the SNR is below 5 dB.
@@ II-B Generative Diffusion Models
$q(\bm{z}_{t}|\bm{x}_{t-1})=\mathcal{N}(\bm{z}_{t};\sqrt{\alpha_{t}}\bm{z}_{t-1},(1-\alpha_{t})\bm{I}),$ (4)
where $\bm{x}_{t}$ is the noisy image at time step $t$, and $\alpha_{t}\in(0,1)$ is a hyperparameter that controls the noise level. Therefore, given a training image $\bm{z}_{0}$, after ${T_{F}}$ steps of forward diffusion, we can directly write the final noisy latent $\bm{z}_{T_{F}}$ through the reparameterization trick as
$\bm{z}_{T_{F}}=\mathcal{N}(\bm{z}_{T_{F}};\sqrt{\alpha_{T_{F}}}\bm{z}_{0},(1-\bar{\alpha}_{T_{F}})\bm{I})$ (5)
where $\overline{\alpha}_{T_{F}}=\prod_{i=1}^{T_{F}}\alpha_{i}$ is the cumulative product of the noise schedule, and $\bm{\epsilon}\sim\mathcal{N}(0,\bm{I})$ is a Gaussian noise. We note that $\alpha_{0}>\alpha_{t}>\cdots>\alpha_{T_{F}}$ is satisfied in the training process to make sure that $\bar{\alpha}_{T_{F}}$ monotonically decreases as $T_{F}$ increases.
@@
$p_{\omega}(\bm{z}_{t-1}|\bm{z}_{t})=\mathcal{N}(\bm{z}_{t-1};\mu_{\omega}(\bm{z}_{t},t),\Sigma_{\omega}(\bm{z}_{t},t)),$ (6)
where $\mu_{\omega}(\bm{z}_{t},t)$ is the predicted mean and $\Sigma_{\omega}(\bm{z}_{t},t)$ is typically a predefined variance. Specifically, the mean $\mu_{\omega}(\bm{z}_{t},t)$ can be further written as
$\mu_{\omega}(\bm{z}_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}(\bm{z}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\overline{\alpha}_{t}}}\epsilon_{\omega}(\bm{z}_{t},t)),$ (7)
where $\epsilon_{\omega}$ is typically modeled as a neural network with U-Net architecture to predict the noise $\epsilon_{\omega}(\bm{z}_{t},t)$ added to the image at each time step. Moreover, it is trained using the following loss function
@@
where $\bm{\epsilon}$ is the ground-truth noise added to the image at time step $t$. After training, the model can generate new images by sampling from the noise distribution $\bm{z}_{T}\sim\mathcal{N}(0,\bm{I})$ and iteratively applying the reverse denoising process in (6) for $T$ steps or performing jump sampling using the DDIM[9] to accelerate the generation process without sacrificing the generation quality, which can be expressed as (9).
@@ III Proposed System
In this section, we introduce the proposed training-free generative SemCom system based on the stable diffusion model. Specifically, we first present the key components of the proposed system, and then we present the details of the semantic encoding and decoding processes. Finally, we provide an analysis of the proposed system.
@@ III-A Overview
As shown in Fig. 1, the proposed system is built upon the components of a stable diffusion model, such as a CLIP text encoder $\text{CLIP}(\cdot)$, an image encoder $\mathcal{E}_{\text{SD}}(\cdot)$, a conditional U-Net model $\epsilon_{\omega}(\cdot,\cdot,\cdot)$, and an image decoder $\mathcal{D}_{\text{SD}(\cdot)}$. Specifically, at the transmitter side, the semantic encoder consists of a CLIP text encoder, an image encoder, and a U-Net model, which is used to extract the semantic features from the input image and generate the channel input signal. At the receiver side, there is also the same CLIP text encoder, U-Net model as the transmitter side, and an image decoder, which is used to reconstruct the image from the received signal. We note that all the components are pre-trained and fixed during the transmission process to align with the training-free design.
@@ III-B Semantic Encoding via DDIM Inversion
To extract the semantic information from the input image, we first use the image encoder to extract the latent feature, i.e.,  $\bm{z}_{0}=\mathcal{E}_{\mathrm{SD}}(\bm{x})$. Unlike the work in [13] that directly transmits the latent feature $\bm{z}_{0}$ as well as the extracted caption, and then adds random noise in the diffusion forward process at the receiver side, we make three key modifications to improve the system performance. Firstly, we propose to use the DDIM inversion to add $T_{F,1}$ steps of deterministic noise to the latent feature $\bm{z}_{0}$ before transmission. This process can be derived by reversing the DDIM sampling process in (9), and can be written as (10), where $\bm{c}$ is a text prompt, $\bm{s}=\text{CLIP}(\bm{c})$ is the corresponding embedding produced by the CLIP text encoder, and $\epsilon_{\omega}(\bm{z}_{t},t,\bm{s})$ is the predicted noise at time step $t$ by the U-Net model conditioned on $\bm{s}$. Thanks to the same U-Net model used in the forward and reverse processes, we can remove the noise more easily than the random noise added in [13]. Secondly, this process is performed at the transmitter side. This is because if we perform the DDIM inversion at the receiver side, the introduced channel noise would distort the latent feature before inversion, making it difficult for the UNet to accurately estimate the noise at each forward timestep, which finally deteriorates the invertibility of DDIM inversion. Thirdly, in the proposed system, we eliminate the need to extract and transmit the image caption, and use a predefined and fixed text prompt such as “High quality, High res, 8K” instead. We note that $T_{F,1}$ forward diffusion steps are performed at the transmitter, and the obtained latent $\bm{z}_{T_{F,1}}$ is normalized by a scaling factor $\gamma=1/\sqrt{\tfrac{1}{2k}||\bm{z}_{T_{F,1}}||_{2}^{2}}$ to satisfy the unit average power constraint [^1] The factor $1/2$ comes from the conversion from real-valued to complex-valued signals.. The normalized latent $\bm{z}=\gamma\bm{z}_{T_{F,1}}$is then mapped into complex-valued symbols for transmission.
@@ III-C Semantic Decoding via DDIM Inversion and Sampling
After receiving the corrupted signal $\bm{y}$, the receiver first converts it into a real-valued vector. Next, the core idea of the proposed system is to set $\hat{\bm{z}}_{T_{F,1}}=\bm{y}$ and continue the DDIM forward process for $T_{F,2}$ steps, so that the noise level of this forward process aligns with the level of channel noise. We refer to the resulting noisy latent feature as $\hat{\bm{z}}_{T_{F}}$, where $T_{F}=T_{F,1}+T_{F,2}$. While these steps can be performed at the transmitter side, it will increase the computation complexity of the transmitter and thus motivate us to split the forward process into two parts, which is similar to the work about split computing. Moreover, we also find that this can effectively mitigate the impact of channel noise.
Next, the receiver can perform the DDIM sampling process in (9) to remove the $T_{F}$ steps of noise added to the latent feature. However, we note that due to the presence of channel noise, the actual noise level in $\bm{z}_{T_{F}}$ is no longer exactly equivalent to that introduced by $T_{F}$ DDIM steps. To address this issue, we propose to define $T_{B}>T_{F}$ as the number of steps to remove the noise added to the latent feature, and set $\tilde{\bm{z}}_{T_{B}}=\bm{z}_{T_{F}}$, and then perform the DDIM sampling process for $T_{B}$ steps to derive the denoised latent feature $\tilde{\bm{z}}_{0}$. Finally, the image decoder is used to reconstruct the image from the latent feature, i.e. $\hat{\bm{x}}=\mathcal{D}_{\mathrm{SD}}(\tilde{\bm{z}}_{0})$. We summarize the pipeline of the proposed system in Algorithm 1.
@@ III-D Analysis
Original Refer to caption MS-SSIM/LPIPS
@@ IV Simulations
In this section, we conduct extensive experiments, including ablation studies, to evaluate the performance of the proposed system.
@@ IV-C Ablation Studies
We also conduct ablation studies to validate the effectiveness of the modifications made in the proposed system. As shown in Table I, we compare the performance of the proposed system with random noise and DDIM inversion noise, where the SNR is set to 5 dB. We also investigate the impact of not transmitting the caption. From this table, we can observe that the proposed system with DDIM inversion noise outperforms the one with random noise in all metrics, which demonstrates the effectiveness of the DDIM inversion used in the proposed system. Moreover, we find that the performance of the proposed system without transmitting the caption is only slightly lower than the version with a losslessly transmitted caption, with a gap of around 3% in FID, which highlights the effectiveness of using a predefined prompt.
@@ V Conclusion
In this paper, we have demonstrated that pretrained GDMs can serve as effective training-free SemCom systems when appropriately adapted. By introducing a two-stage forward diffusion strategy and analyzing the impact of channel noise on the latent distribution, we have revealed how diffusion-based semantic encoding and denoising processes can align with noisy transmission environments. Through extensive experiments, we have validated that the proposed system has achieved superior performance across diverse evaluation metrics.

### v2

@@ Abstract
Semantic communication, deep joint source-channel coding, diffusion models, image transmission.
@@ I Introduction
To eliminate the dependency on training, recent studies have explored training-free generative SemCom frameworks. Specifically, the authors in [12] introduced a channel-aware GAN inversion method for semantic encoding and employed the same GAN generator for decoding, thus avoiding any encoder–decoder training on the communication task. Meanwhile, the works in [13, 14] proposed to use a pretrained image captioner to extract textual descriptions from source images. These captions, along with auxiliary visual features such as edge maps or latent vectors, are transmitted to the receiver, where the caption serves as a conditioning prompt to guide a GDM in reconstructing the original image. However, these training-free approaches lack robustness to channel noise, as the transmitted semantic conditions, such as captions or latent features, are easily corrupted during transmission, leading to degraded reconstruction quality.
To overcome these limitations, we propose a fully training-free generative SemCom framework that leverages publicly available pretrained GDMs. Specifically, we design a semantic encoding and decoding method based on the sampling and inversion processes of the denoising diffusion implicit model (DDIM) [9], which introduces a two-stage forward diffusion strategy split between the transmitter and receiver to enhance robustness against channel noise. In addition, we optimize the number of sampling steps at the receiver to match the total noise level introduced by the channel noise. We further provide a brief analysis of this design, how channel noise shifts the latent distribution away from the latent distribution of GDM in ideal conditions, and give insights into how the proposed system mitigates this mismatch. Simulations on the Kodak dataset are conducted to demonstrate the superiority of the proposed system across a wide range of perceptual and distortion metrics. In particular, our method achieves over 50% performance gains in Fréchet Inception distance (FID) compared to baselines when the SNR is below 5 dB.
@@ II-B Generative Diffusion Models
$q(\bm{z}_{t}|\bm{x}_{t-1})=\mathcal{N}(\sqrt{\alpha_{t}}\bm{z}_{t-1},(1-\alpha_{t})\bm{I}),$ (4)
where $\bm{x}_{t}$ is the noisy image at time step $t$, and $\alpha_{t}\in(0,1)$ is a hyperparameter that schedules the noise level. Therefore, given a training image $\bm{z}_{0}$, after ${T_{F}}$ steps of forward diffusion, we can directly write the final noisy latent $\bm{z}_{T_{F}}$ through the reparameterization trick as
$\bm{z}_{T_{F}}\sim\mathcal{N}(\sqrt{\alpha_{T_{F}}}\bm{z}_{0},(1-\bar{\alpha}_{T_{F}})\bm{I}),$ (5)
where $\overline{\alpha}_{T_{F}}=\prod_{i=1}^{T_{F}}\alpha_{i}$ is the cumulative product of the noise schedule terms, and $\bm{\epsilon}\sim\mathcal{N}(0,\bm{I})$ is a Gaussian noise. We note that $\alpha_{0}>\alpha_{t}>\cdots>\alpha_{T_{F}}$ is satisfied in the training process to make sure that $\bar{\alpha}_{T_{F}}$ monotonically decreases as $T_{F}$ increases.
@@
$p_{\omega}(\bm{z}_{t-1}|\bm{z}_{t})=\mathcal{N}(\mu_{\omega}(\bm{z}_{t},t),\Sigma_{\omega}(\bm{z}_{t},t)),$ (6)
where $\Sigma_{\omega}(\bm{z}_{t},t)$ is a predefined variance and $\mu_{\omega}(\bm{z}_{t},t)$ is the predicted mean, given as
$\mu_{\omega}(\bm{z}_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}(\bm{z}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\overline{\alpha}_{t}}}\epsilon_{\omega}(\bm{z}_{t},t)).$ (7)
The denoising term $\epsilon_{\omega}$ is typically modeled as a neural network with U-Net architecture to predict the noise $\epsilon_{\omega}(\bm{z}_{t},t)$ to be subtracted from the image at each time step. This neural network is trained using the following loss function:
@@
where $\bm{\epsilon}$ is the ground-truth noise to be subtracted from the image at time step $t$. After training, the model can generate new images by sampling from the noise distribution $\bm{z}_{T}\sim\mathcal{N}(0,\bm{I})$ and iteratively applying the reverse denoising process in (6) for $T$ steps or performing jump sampling using the DDIM [9] to accelerate the generation without compromising quality, which can be expressed as (9) in the next page.
@@ III Proposed System
In this section, we introduce the proposed training-free generative SemCom system built upon the stable diffusion model, a DDIM-based latent diffusion model. We first present the key components of the proposed system, and then elaborate the details of the semantic encoding and decoding processes. Finally, we provide an analysis of the proposed system.
@@ III-A Overview
Following the stable diffusion model, our proposed generative SemCom system comprises a CLIP text encoder $\text{CLIP}(\cdot)$, an image encoder $\mathcal{E}_{\text{SD}}(\cdot)$, a conditional U-Net model $\epsilon_{\omega}(\cdot,\cdot,\cdot)$, and an image decoder $\mathcal{D}_{\text{SD}(\cdot)}$, as Fig. Fig. 1 illustrates. At the transmitter side, the semantic encoder consists of a CLIP text encoder, an image encoder, and a U-Net model, which is used to extract the semantic features from the input image and generate the channel input signal. At the receiver side, there is also the same CLIP text encoder, U-Net model as the transmitter side, and an image decoder, which is used to reconstruct the image from the received signal. We note that all the components are pre-trained and fixed during the transmission process to align with the training-free design.
@@ III-B Semantic Encoding via DDIM Inversion
To extract the semantic information from the input image, we first use the image encoder to extract the latent feature, i.e.,  $\bm{z}_{0}=\mathcal{E}_{\mathrm{SD}}(\bm{x})$. Unlike the work in [13] that directly transmits the latent feature $\bm{z}_{0}$ as well as the extracted caption, and then adds random noise in the diffusion forward process at the receiver side, we make three key modifications to improve the system performance as follows.
  1. 1.
    We propose to use the DDIM inversion to add $T_{F,1}$ steps of deterministic noise to the latent feature $\bm{z}_{0}$ before transmission. This process can be derived by reversing the DDIM sampling process in (9), and can be written as (10) at the top of this page. The term $\bm{c}$ is a text prompt, $\bm{s}=\text{CLIP}(\bm{c})$ is the corresponding embedding produced by the CLIP text encoder, and $\epsilon_{\omega}(\bm{z}_{t},t,\bm{s})$ is the predicted noise at time step $t$ by the U-Net model conditioned on $\bm{s}$. Thanks to the same U-Net model used in the forward and reverse processes, we can remove the noise more easily than the random noise added in [13].
  2. 2.
    The inversion process is performed at the transmitter side. This is because if we perform the DDIM inversion at the receiver side, the noisy channel would distort the latent feature before inversion. This makes the UNet difficult to accurately estimate the noise at each forward timestep, deteriorating the invertibility of DDIM inversion.
  3. 3.
    Instead of extracting and transmitting the image caption, we use a predefined text prompt such as “High quality, High res, 8K,” which our experiments found to be more effective in improving reconstruction quality. We note that $T_{F,1}$ forward diffusion steps are performed at the transmitter, and the obtained latent $\bm{z}_{T_{F,1}}$ is normalized by a scaling factor $\gamma=1/\sqrt{\tfrac{1}{2k}||\bm{z}_{T_{F,1}}||_{2}^{2}}$ to satisfy the unit average power constraint [^1] The factor $1/2$ comes from the conversion from real-valued to complex-valued signals.. The normalized latent $\bm{z}=\gamma\bm{z}_{T_{F,1}}$is then mapped into complex-valued symbols for transmission.
@@ III-C Semantic Decoding via DDIM Inversion and Sampling
Upon receiving the noisy signal $\bm{y}$, the receiver first converts it into a real-valued vector. Next, the core idea of the proposed system is to set $\hat{\bm{z}}_{T_{F,1}}=\bm{y}$ and continue the DDIM forward process for $T_{F,2}$ steps, so that the noise level of this forward process aligns with the level of channel noise. We refer to the resulting noisy latent feature as $\hat{\bm{z}}_{T_{F}}$, where $T_{F}=T_{F,1}+T_{F,2}$. The rationale behind this forward process splitting is to avoid channel noise prediction at the transmitter by performing part of the forward process after channel perturbation at the receiver. Given this split architecture, we will provide a design guideline for the forward and backward processes in the next section. Note that when the receiver’s computational capability is severely limited, a non-split architecture may be preferable. Optimizing such split model partitioning is an interesting direction for future study.
Next, the receiver performs the DDIM sampling process in (9) to remove the $T_{F}$ steps of noise added to the latent feature. However, we note that due to the presence of channel noise, the actual noise level in $\bm{z}_{T_{F}}$ is no longer exactly equivalent to that introduced by $T_{F}$ DDIM steps. To address this issue, we propose to define $T_{B}>T_{F}$ as the number of steps to remove the noise added to the latent feature, and set $\tilde{\bm{z}}_{T_{B}}=\bm{z}_{T_{F}}$, and then perform the DDIM sampling process for $T_{B}$ steps to derive the denoised latent feature $\tilde{\bm{z}}_{0}$. Finally, the image decoder is used to reconstruct the image from the latent feature, i.e. $\hat{\bm{x}}=\mathcal{D}_{\mathrm{SD}}(\tilde{\bm{z}}_{0})$. We summarize the pipeline of the proposed system in Algorithm 1.
@@ III-D Analysis
Original Refer to caption PSNR/LPIPS
@@ IV-C Ablation Studies
We additionally conduct ablation studies to validate the effectiveness of the modifications made in the proposed system. As shown in Table I, we compare the performance of the proposed system with random noise and DDIM inversion noise, where the SNR is set to 5 dB. We also investigate the impact of not transmitting the caption. From this table, we can observe that the proposed system with DDIM inversion noise outperforms the one with random noise in all metrics, which demonstrates the effectiveness of the DDIM inversion used in the proposed system. Moreover, we find that the performance of the proposed system without transmitting the caption is only slightly lower than the version with a losslessly transmitted caption, with a gap of around 3% in FID, which highlights the effectiveness of using a predefined prompt.
@@ V Conclusion
In this paper, we have demonstrated that pretrained GDMs can serve as effective training-free JSCC for SemCom systems when appropriately adapted. By introducing a two-stage forward diffusion strategy and analyzing the impact of channel noise on the latent distribution, we have revealed how diffusion-based semantic encoding and denoising processes can align with noisy channel environments. Through extensive experiments, we have validated that the proposed system outperforms existing SemCom frameworks across diverse evaluation metrics.